# A Learnable Prior Improves Inverse Tumor Growth Modeling

### v1

@@ Abstract
Figure 1: Graphical Abstract: First, a DL network predicts a prior of tumor model parameters (orange) based on imaging data. Second, this prior is used as an initialization for the subsequent sampling-based optimization strategy (blue dots) to predict the final tumor model parameters, leading to the full modeling of tumor concentration.
@@ 1.1 Glioblastoma
Brain tumors are a significant public health concern, with high morbidity and mortality rates. Glioblastoma is the most common primary malignant brain tumor with a median survival time of only 14-17 months [1] and a yearly incidence rate of 3.53 per 100,000 [2]. Despite recent advances in understanding the biology of this disease and prolonged survival with modern combinations of radio- and chemotherapy [3], there is still a tremendous need for improving the treatment of these patients. A central challenge is the tendency of glioblastoma cells to diffusely spread from the primary tumor into the surrounding brain.
1.2 Radiotherapy
Radiotherapy represents a central pillar in brain tumor treatment, as it can be used as a primary treatment, adjuvant treatment, or palliative treatment to improve quality of life. The goal of radiation therapy for brain tumors is striking a delicate balance between delivering an effective dose of radiation to the tumor and minimizing damage to surrounding healthy brain tissue.
@@ 1.2 Radiotherapy
Current glioblastoma radiation therapy planning predominantly relies on MR and CT imaging, with the clinical target volume (CTV) defined as a uniform 15 mm margin around the resection cavity plus T1 contrast-enhancing tumor (GTV) to account for the diffuse infiltration of tumor cells into the surrounding brain [4]. While this approach aims to address the diffuse spread, it lacks true patient individualization due to its uniformity. Imaging-based individualization of radiotherapy promises to meet the need for more effective therapy.
1.3 Tumor Modeling
@@
Radiotherapy planning is limited to the visible boundaries of the tumor on imaging modalities, which is known to only show the “tip of the iceberg”. Physical modeling has been explored to predict tumor cell concentrations based on growth patterns to overcome this limitation and uncover the full extent of the tumor. These models aim to incorporate tumor pathophysiology through the spatial and temporal dynamics of the tumor. Various models are used, ranging from agent-based cellular automaton over multi-scale models to grid-based continuous models, which are applied in this paper [5, 6, 7, 8, 9].
@@
Traditionally, various sampling-based approaches are used to solve the inverse tumor problem, i.e., to estimate the growth model parameters best describing the observed tumor. Simply put, these methods run many forward simulations in an optimized way to converge toward the most likely set of parameters or towards an entire posterior probability distribution of the parameters [10]. For instance, Bayesian inference executed by Markov-chain Monte Carlo can lead to high-quality model fits [11]. With typical forward solvers using the Lattice Boltzmann methods and multi-scope adapted grid methods, a runtime of about 1-3 min per forward run can be achieved for simple tumor models [10, 11]. However, multiple hundred up to a few thousand samples must be simulated for a reasonable inference result. Thus, although sampling strategies are very precise and, in principle, attractive for clinical use given their convergence robustness, they have a strongly growing complexity with the number of parameters relevant to advanced models. Therefore, these methods are prohibitively costly to calculate, forming a significant hurdle for their routine clinical use.
@@
Recently, neural networks have been employed to estimate parameters for brain tumor models [12, 13, 14] or tumor growth volumes directly [15]. Convolutional neural networks (CNNs) are designed to analyze images and identify patterns, making them well-suited for analyzing medical images and estimating growth model parameters. Applying CNNs for brain tumor modeling involves training the network to recognize specific features within the MR images associated with tumor growth and progression. After training, these networks can solve the inverse problem of predicting tumor model parameters from an imaging observation to drive a deterministic physics simulation. Despite the fast inference time, the main drawback of a pure DL approach lies in its inflexibility towards modal adaptations and the lack of error guarantees, which can lead to incorrect parameter estimation [16]. This lack of robustness represents a significant challenge for their clinical translation.
@@
1.4 Contribution
@@ 2.1 Framework Overview
Our goal is to predict the tumor concentration for each voxel, which is not visible in conventional imaging modalities, based on physically constrained modeling (Figure 2). As input to our model, contrast-enhanced T1 (T1c) and FLAIR MRI scans are used, highlighting higher and lower tumor concentrations, respectively (Figure 2 a). In a preprocessing step, the tumor regions are segmented in both scans into the subregions contrast-enhancing tumor (CET) and edema, using BraTS Toolkit [18]. The tissue segmentation is generated based on the T1c scan.
The tumor and tissue segmentations are fed into a neural network to predict the physical tumor model parameters (Figure 2 b). The premise is that the network learns the growth dynamics based on the underlying tissue. Therefore, an overlay of both tumor segmentations is fed into the network as a single channel, whereas the other input channels contain the patient-specific tissue information (WM, GM, CSF) to facilitate patient-specific modeling. The network output consists of scalar parameters that determine the physical process of tumor development. From those parameters, it is possible to simulate the growth process of the tumor within the patient’s brain.
@@ 2.2 Tumor Growth Model
The first term describes the diffusion process with diffusion coefficient $D$, and the second term describes the growth process over time $t$. In the case of the applied Fisher-Kolmogorov equation, a logistic growth process with a growth parameter $\rho$ is assumed.
@@ 2.3 Tumor Simulation
For the simulation, three types of brain tissue are distinguished: white matter (WM), gray matter (GM), and cerebrospinal fluid (CSF). For each voxel $i$ the probability $p$ of all tissues adds up to one: $p_{{}_{WM}}+p_{{}_{GM}}+p_{{}_{CSF}}=1$.
@@
Additional parameters are the tumor origin $\{x,y,z\}$ and the total time of tumor growth $T$. Thus, the original set of parameters for the forward solver is $\theta_{orig}=\{x,y,z,\rho,D_{{}_{WM}},T\}$. As only one time-point is realistically available for radiotherapy planning, $\theta_{orig}$ is ambiguous: for example, a faster-growing tumor over a shorter time and a slower-growing tumor over a longer time are indistinguishable based on a single time point. Therefore we introduce end-time independent parameters $\mu_{D}=\sqrt{D_{{}_{WM}}T}$ [cm] and $\mu_{\rho}=\sqrt{\rho T}$ [unitless] leading to the reduced parameter set: $\theta=\{x,y,z,\mu_{D},\mu_{\rho}\}$. The 3D tumor cell density $c$ is a function of these parameters. For simulation, the GliomaSolver [11] software is used.
2.4 Preprocessing
For the network training, 30,000 simulations are created. To allow the network to learn tumor dynamics in a large variety of patient brain structures, 140 real patient scans are utilized, taken from the BraTS dataset [19]. For testing, 180 simulations are conducted in another 140 BraTS patient anatomies with random parameters. The tissue is segmented into WM, GM, and CSF using the S3 software [11].
@@ 2.5.1 Network Training
The network encoder (Figure 2 b) is based on a residual neural network (ResNet) [20]. In contrast to former experiments [13], the network input consists of four channels with dimensions $128\times 128\times 128$, concatenating the three patient-specific tissue masks and the tumor segmentation map. The network outputs are the five model parameters $\theta$ optimizing an $L_{2}$ loss. The simulation parameters are sampled based on clinical experience in the same way as introduced in the Learn-Morph-Infer (LMI) method [13].
@@ 2.5.2 Network Ensemble
An ensemble of ten networks is introduced to increase the model robustness towards multiple edema and CET thresholds. The last layer of the network ensemble is fine-tuned for a certain threshold ($c^{th}_{Edema},c^{th}_{CET}$) in the ground truth training data. The premise is that the ensemble mean is more stable toward single network failure, and the ensemble standard deviation can quantify the uncertainty of the prediction.
@@ 2.6.1 Prior
A Gaussian distribution around the DL estimated parameters is defined as prior. The prior distribution is based on the ensemble mean ($\bar{\theta}^{i}_{dl}$) and its standard deviation ($\sigma^{i}_{dl}$) for each parameter $i\in\theta$. It is assumed that the individual parameters follow independent Gaussian distributions.
@@ 2.6.4 Sampling Algorithm
To ensure an equitable comparison of runtime between sampling methods with and without a prior, we implement a specific initialization protocol for the absence of a prior. The tumor origin is initialized at the center of mass of the FLAIR segmentation. The diffusion coefficient $\mu_{D}$ and $\mu_{rho}$ are initialized at its lower limit, given that the forward simulation runtime increases linearly with $\mu_{D}$. This initialization results in a sampling strategy that converges towards optimal parameters while commencing with a small tumor and subsequently increasing in size. As smaller tumors facilitate expedited simulation time, this initialization improves the entire sampling runtime. For this comparative analysis, a subset of 13 patients was randomly selected for cost reasons. We sampled until a convergence up to 99 % of the maximum posterior up to a total of 2000 samples for the case without prior and 600 samples for the case with DL-Prior.
@@ 3.1.2 Estimated Parameter Comparison
The estimation of the tumor origin is generally precise, with an average deviation of 3.7 mm to the ground truth for the single network (3.6 mm for the ensemble mean). This is twice as good as the LMI approach, with a mean displacement of 8.7 mm. As a baseline, the edema segmentation’s center of mass is included as a rough estimation. The center of mass mean displacement is 15.9 mm away from the ground truth, demonstrating the superiority of our method.
@@
Table 1: The tumor model parameter mean displacement (MD) and the root-means-squared deviation (RMSD) from the simulation ground truth are shown for different methods. Additionally, the distance from the center of mass to the proposed tumor origin is included.
Metric Parameter DL-Prior + Sampling DL DL Ensemble LMI Edema Center of Mass
Origin [mm] 3.51 3.72 3.57 8.72 15.9
MD $\mu_{D}$ [mm] -0.30 -0.24 -0.25 1.65
$\mu_{\rho}$ 0.01 -0.02 0.05 0.01
Origin [mm] 4.38 4.55 4.37 10.30 18.65
RMSD $\mu_{D}$ [mm] 1.05 0.84 0.82 2.53
$\mu_{\rho}$ 0.40 0.33 0.33 0.74
@@ 3.1.3 Estimated Tumor Concentration
Table 2: The mean absolute (MAE) and mean squared error (MSE) are compared between the predicted tumor concentration and the ground truth for different algorithms.
MAE [$10^{-4}$] MSE [$10^{-5}$]
DL-Prior + Sampling 16 $\pm$ 2 30$\pm$ 7
DL 24 $\pm$ 2 68 $\pm$ 10
DL Ensemble 27 $\pm$ 3 72 $\pm$ 11
LMI 79 ±$\pm 6$ 401$\pm$ 40

### v2

@@ Abstract
J. Weidner, I. Ezhov, L. Feiner und L. Lux are with the Dept. of Computer Science and the Center for Translational Cancer Research, TranslaTUM, Technical University of Munich, Germany (e-mail: j.weidner@tum.de, ivan.ezhov@tum.de, leo.feiner@tum.de, laurin.lux@tum.de).
B. Wiestler and M. Metz are with the Dept. of Neuroradiology, School of Medicine and Health, Technical University of Munich, Germany (e-mail: b.wiestler@tum.de, marie.metz@tum.de)
S. Litvinov and S. Kaltenbach are with the School of Engineering and Applied Sciences, Harvard University, MA, USA (e-mail: slitvinov@seas.harvard.edu, skaltenbach@seas.harvard.edu)
Michal Balcerak and Björn Menze are with the Dept. of Quantitative Biomedicine, University of Zurich, Zurich, Switzerland (e-mail: michal.balcerak@uzh.ch, bjoern.menze@uzh.ch)
F. Kofler is with Helmholtz AI, Helmholtz Zentrum Munich, Germany and with the Center for Translational Cancer Research, TranslaTUM, Technical University of Munich (e-mail: florian.kofler@helmholtz-munich.de)
J. Lipkova is with the Department of Pathology, University of California Irvine, CA, USA (e-mail: jlipkova@hs.uci.edu)
J. Latz is with the Department of Mathematics, University of Manchester, UK (e-mail: jonas.latz@manchester.ac.uk)
D. Rueckert is with the Department of Computing, Biomedical Image Analysis Group, Imperial College London, UK and with the Klinikum rechts der Isar, Technical University Munich, Germany (e-mail: daniel.rueckert@tum.de)
Bjoern Menze and Benedikt Wiestler share senior authorship.
This work was supported in part by the DFG (Deutsche Forschungsgemeinschaft) under Grant 504320104 and by the European High Performance Computing Joint Undertaking (EuroHPC) Grant DCoMEX (956201-H2020-JTI-EuroHPC-2019-1)
Benedikt Wiestler is supported by the NIH (R01CA269948).
Bjoern Menze and Michal Balcerak acknowledge support by the Helmut-Horten-Foundation.
@@ 1 Introduction
Glioblastoma is the most common primary malignant brain tumor with a median survival time of only 14-17 months [1] and a yearly incidence rate of 3.53 per 100,000 [2]. Despite recent advances in understanding the biology of this disease and prolonged survival with modern combinations of radio- and chemotherapy [3], there is still a tremendous need for improving the treatment of these patients. A central challenge is the tendency of glioblastoma cells to diffusely spread from the primary tumor into the surrounding brain.
@@
Radiotherapy represents a central pillar in brain tumor treatment. The goal of radiation therapy for brain tumors is striking a delicate balance between delivering an effective dose of radiation to the tumor and minimizing damage to surrounding healthy brain tissue. Current glioblastoma radiation therapy planning predominantly relies on MR and CT imaging, with the clinical target volume defined as a uniform 15 mm margin around the resection cavity plus T1 contrast-enhancing tumor [4]. While this approach aims to address the diffuse spread, it lacks true patient individualization due to its uniformity. Imaging-based individualization of radiotherapy promises to meet the need for more effective therapy.
@@
Radiotherapy planning is limited to the visible boundaries of the tumor on imaging modalities, which is known to only show the “tip of the iceberg” [4]. Physical modeling has been explored to predict tumor cell concentrations based on growth patterns to overcome this limitation and uncover the full extent of the tumor. These models aim to incorporate tumor pathophysiology through the spatial and temporal dynamics of the tumor. In the literature, various models are used, ranging from agent-based cellular automaton over multi-scale models to grid-based continuous models, which are applied in this paper [5, 6, 7, 8, 9].
@@
Traditionally, various sampling-based approaches are used to solve the inverse tumor problem, i.e., to estimate the growth model parameters best describing the observed tumor. Simply put, these methods run many forward tumor growth simulations [10], called samples, in an optimized way to converge toward the most likely set of parameters or towards an entire posterior probability distribution of the parameters [11]. For instance, Bayesian inference executed by Markov-chain Monte Carlo can lead to high-quality model fits [12]. With typical forward solvers using the Lattice Boltzmann methods and multi-scope adapted grid methods, a runtime of about 1-3 min per forward run can be achieved for simple tumor models [11, 12]. However, multiple hundred up to a few thousand samples must be simulated for a reasonable inference result. Thus, although sampling strategies are very precise and, in principle, attractive for clinical use given their convergence robustness, they have a strongly growing complexity with the number of parameters relevant to advanced models. Therefore, these methods are prohibitively costly to calculate, running multiple days per patient [12], forming a significant hurdle for their routine clinical use.
@@
Recently, neural networks have been employed to estimate parameters for brain tumor models [13, 14, 15, 16] or tumor growth volumes directly [17]. Convolutional neural networks (CNNs) are designed to analyze images and identify patterns, making them well-suited for analyzing medical images and estimating growth model parameters. Applying CNNs for brain tumor modeling involves training the network to recognize specific features within the MR images associated with tumor growth and progression. After training, these networks can solve the inverse problem of predicting tumor model parameters from an imaging observation to drive a deterministic physics simulation. Despite the fast inference time of less than 5 minutes, the main drawback of a pure DL approach lies in its inflexibility towards modal adaptations and the lack of error guarantees, which can lead to incorrect parameter estimation [18]. This lack of robustness represents a significant challenge for their clinical translation.
@@
The successful integration of deep learning with traditional methods for diverse inverse problems across various fields [20, 21, 22] motivated us to explore this approach for brain tumors.
Contributions
@@ 2.1 Framework Overview
Our goal is to predict the tumor concentration for each voxel, which is not visible in conventional imaging modalities, based on physically constrained modeling (Figure 1). As input to our model, contrast-enhanced T1 (T1c) and FLAIR MRI scans are used, highlighting higher and lower tumor concentrations, respectively (Figure 1 a). In a preprocessing step, the tumor regions are segmented in both scans into the subregions contrast-enhancing tumor (CET) and edema, using BraTS Toolkit [23]. The tissue segmentation is generated based on the T1c scan using the S3 software [12].
The tumor and tissue segmentations are fed into a neural network to predict the physical tumor model parameters (Figure 1 b). The premise is that the network learns the growth dynamics based on the underlying tissue. Therefore, an overlay of both tumor segmentations is fed into the network as a single channel(edema: 0.33, CET:0.66), whereas the other input channels contain the patient-specific tissue information: white matter (WM), gray matter (GM) and cerebrospinal fluid (CSF) to facilitate patient-specific modeling. The tissue information in each channel is continuous and can be interpreted as the ratio of the given tissue type per voxel. The network output consists of scalar parameters that determine the physical process of tumor development. From those parameters, it is possible to simulate the growth process of the tumor within the patient’s brain.
@@ 2.2 Tumor Growth Model
The first term describes the diffusion process with diffusion coefficient $D$, and the second term describes the growth process over time $t$. In the case of the applied Fisher-Kolmogorov equation [24], a logistic growth process with a growth parameter $\rho$ is assumed.
@@ 2.3 Tumor Simulation
For the simulation, three types of brain tissue are distinguished: WM, GM and CSF. For each voxel $i$ the probability $p$ of all tissues adds up to one: $p_{{}_{WM}}+p_{{}_{GM}}+p_{{}_{CSF}}=1$.
@@
Additional parameters are the tumor origin $\{x,y,z\}$ and the total time of tumor growth $T$. We select a single tumor origin as the starting point of the simulation. This simulated origin cannot necessarily be directly associated with the actual tumor-originating cell in medical cases. Thus, the original set of parameters for the forward solver is $\theta_{orig}=\{x,y,z,\rho,D_{{}_{WM}},T\}$. As only one time-point is realistically available for radiotherapy planning, $\theta_{orig}$ is ambiguous: for example, a faster-growing tumor over a shorter time and a slower-growing tumor over a longer time are indistinguishable based on a single time point. Therefore we introduce end-time independent parameters $\mu_{D}=\sqrt{D_{{}_{WM}}T}$ [cm] and $\mu_{\rho}=\sqrt{\rho T}$ [unitless] leading to the reduced parameter set: $\theta=\{x,y,z,\mu_{D},\mu_{\rho}\}$. The 3D tumor cell density $c$ is a function of these parameters. For simulation, the GliomaSolver [12] software is used.
2.4 Data
For the network training, 30,000 simulations are created using 140 real patient tissues from the BraTS dataset [25] to simulate synthetic tumors. By explicitly incorporating tissues from BraTS patients, whose anatomy is already altered by tumors, we introduce pathological anatomy variations that the network is likely to encounter in real patient data. This approach aims to make the model invariant to tumor-induced anatomical differences and typical tissue deformations, thereby enhancing its ability to generalize to diverse patient data. The dataset is split into 28000 training samples and 2000 validation samples.
For the additional test set for extensive validation of all methods, including costly additional sampling, 180 simulations are conducted in another 140 BraTS patient anatomies with random parameters.
@@ 2.5.1 Network Training
The network encoder (Figure 1 b) is based on a residual neural network (ResNet) [26]. In contrast to former experiments [14], the network input consists of four channels with dimensions $128\times 128\times 128$, concatenating the three patient-specific tissue masks and the tumor segmentation map. All other parameters and conditions are maintained as in the prior experiments [14]. The network outputs are the five model parameters $\theta$ optimizing an $L_{2}$ loss. The simulation parameters are sampled based on clinical experience in the same way as introduced in the Learn-Morph-Infer (LMI) method [14].
@@
The training was conducted on an NVIDIA Quadro RTX 8000 with 50GB RAM. We trained for 85 hours over 50 epochs.
@@ 2.5.2 Network Ensemble
An ensemble of ten networks is introduced to increase the model robustness towards multiple edema and CET thresholds. The last layer of the network ensemble is fine-tuned for a certain threshold ($c^{th}_{Edema},c^{th}_{CET}$) in the ground truth training data. The premise is that the ensemble mean is more stable toward single network failure, and the ensemble standard deviation can quantify the uncertainty of the prediction. Technically, we fixed all weights except the last linear layer and retrained it on different thresholds. A Gaussian prior is generated from this equally weighted ensemble (see section 2.6.1).
@@ 2.6.1 Prior
A Gaussian distribution around the DL-ensemble estimated parameters is defined as prior. The prior distribution is based on the ensemble mean ($\bar{\theta}^{i}_{dl}$) and its standard deviation ($\sigma^{i}_{dl}$) for each parameter $i\in\theta$. It is assumed that the individual parameters follow independent Gaussian distributions.
@@ 2.6.4 Sampling Algorithm
To ensure an equitable comparison of runtime between sampling methods with and without a prior, we implement a specific initialization protocol for the absence of a prior, the naive sampling. The tumor origin is initialized at the center of mass of the FLAIR segmentation. The diffusion coefficient $\mu_{D}$ and $\mu_{rho}$ are initialized at its lower limit, given that the forward simulation runtime increases linearly with $\mu_{D}$. This initialization results in a sampling strategy that converges towards optimal parameters while commencing with a small tumor and subsequently increasing in size. As smaller tumors facilitate expedited simulation time, this initialization improves the entire sampling runtime. For this comparative analysis, a subset of 13 patients was randomly selected for cost reasons. We sampled until a convergence up to 99 % of the maximum posterior up to a total of 2000 samples for the case without prior and 600 samples for the case with DL-Prior. All simulations were conducted on an AMD EPYC 7313 16-core CPU. We published our code on GitHub222github.com/jonasw247/a-learnable-prior-improves-inverse-tumor-growth-modeling.
@@ 2.7 Evaluation
To ensure statistical robustness, we have highlighted the significant results by marking them with an asterisk (*). We set the threshold for significance to a p-value of 1%. If not stated otherwise, we compared the different methods to our proposed method, ”DL-Prior + Sampling”. As we cannot assume Gaussian distributions and as we analyze paired data, we used the Wilcoxon signed-rank test.
@@ 3.1.2 Estimated Parameter Comparison
The estimation of the tumor origin is generally precise, with an average deviation of 3.7 mm to the ground truth for the single network (3.6 mm for the ensemble mean). This is twice as good as the LMI approach, with a mean displacement of 8.7 mm. As a baseline, the edema segmentation’s center of mass is included as a rough estimation. The CET center of mass mean displacement is 12.7 mm away from the ground truth, demonstrating the superiority of our method.
@@
Table 1: The tumor model parameter mean displacement (MD) and the root-means-squared deviation (RMSD) from the simulation ground truth are shown for different methods. MD is a criterion for the systematic deviation, the shift between the means. RMSD is a measure of the spread of differences in distribution and, thus, determines the quality of the fit. The distance from the edema and CET center of mass to the ground truth tumor origin is included. A significant improvement in origin estimation of DL-Prior + sampling is found over the core center of mass. For $\mu_{D}$ and $\mu_{\rho}$, subsequent sampling results in worse parameter results. This highlights that the problem is ill-posed, meaning that finding a better fit to the observed tumor data can result in deviations or instability in the estimated parameter values (see also Table 2)
Metric Parameter DL-Prior + Sampling DL DL Ensemble LMI Edema Center of Mass CET Center of Mass
Origin [mm] 3.51 $\pm$ 0.19 3.72 $\pm$ 0.20 3.57$\pm$ 0.19 8.72 $\pm$ 0.27 * 15.87 $\pm$ 0.73 * 12.66 $\pm$ 0.56 *
MD $\mu_{D}$ [mm] -0.30$\pm$ 0.08 -0.24$\pm$ 0.06 -0.25$\pm$ 0.06 1.66$\pm$ 0.15 *
$\mu_{\rho}$ 0.01 $\pm$ 0.02 -0.02 $\pm$ 0.02 0.05 $\pm$ 0.02 0.01 $\pm$ 0.06
Origin [mm] 4.38 $\pm$ 0.24 4.56 $\pm$ 0.24 4.37 $\pm$ 0.24 10.30 $\pm$ 0.31 * 18.64 $\pm$ 0.88 * 14.9$\pm$ 0.96 *
RMSD $\mu_{D}$ [mm] 1.05 $\pm$ 0.10 0.84 $\pm$ 0.06 0.82 $\pm$ 0.06 2.56 $\pm$ 0.13 *
$\mu_{\rho}$ 0.40 $\pm$ 0.06 0.33 $\pm$ 0.03 0.33 $\pm$ 0.03 0.74 $\pm$ 0.07
@@ 3.1.3 Estimated Tumor Concentration
Table 2: The mean absolute (MAE) and mean squared error (MSE) are compared between the predicted tumor concentration and the ground truth for different algorithms. A significant improvement of DL-Prior + sampling over the second-best method is demonstrated.
MAE [$10^{-4}$] MSE [$10^{-5}$] Dice 0.5
DL-Prior + Sampling 16 $\pm$ 2 30 $\pm$ 7 0.93 $\pm$ 0.01
DL 24 $\pm$ 2 * 68 $\pm$ 10 * 0.88 $\pm$ 0.01 *
DL Ensemble 27 $\pm$ 3 * 72 $\pm$ 11 * 0.89 $\pm$ 0.01*
LMI 79 ±$\pm 6$ 401$\pm$ 40 * 0.75 $\pm$ 0.01 *